# Flow Graph-Based Classification of Defects4J Faults

### v1

@@ IV Dataset Statistics
Section III shows the frequency of each fault class, in terms of the number of faults it is assigned to, over the 488 classified Defects4J faults. We see that definition fault (data-flow) is the most common fault class for each project. On the other hand, unconditional statement order fault (control-flow) is the least common class. Overall, the total control-flow classifications clearly outweigh the total data-flow classifications, despite individual control-flow classes being less frequently assigned.
@@
We also see the distribution of the number of fault classes by the average number assigned per fault. All faults are assigned, on average, around two fault classes, with Time displaying the highest average ($2.63$). Furthermore, standard deviations of around $1.0$ show that faults are generally assigned between one and three fault classes. Figure 0(a) shows the co-assignments between our fault classes. Each cell $i,j$ shows the percentage of faults assigned fault class $i$ that are also assigned fault class $j$. There is high agreement between several classes, especially with definition fault. Due to the discussed distribution, faults are often assigned more than one class, and definition fault is the most frequent class.
Because the fault classes are not disjoint, we consider disjoint fault types, where a fault is assigned only control-flow classes (pure CF), only data-flow classes (pure DF), or both (CF/DF). The majority of faults display a mixture of control- and data-flow characteristics.
@@
Sobreira et al. [2] investigate 395 Defects4J faults; however, they focus on the repair actions and patterns, rather than fault types. The repair actions are very fine-grained, since they distinguish between addition, removal, and modification in nearly all syntactic categories. This leads to 28 different classes. The nine repair patterns are more abstract, and thus closer to our classification. However, some repair pattern categories are not useful for our purposes. For example, single line does not give much information about the fault itself—the fault is fixed in one line, but this is not the only possible fix. Figure 0(b) shows the co-assignments of seven repair classes from [2], which show fewer agreements than our own. However, we argue that certain of their repair classifications are biased by the fix, and do not accurately represent the underlying fault. Consider LABEL:listing:lang7 and LABEL:listing:math46:
- if (str.startsWith("--")) {
- return null;
Listing 11: Conditional block repair—removal (Lang-$7$[2]
- return isZero ? NaN : INF;
+ return NaN;
Listing 12: Unwraps-from repair—if-else (Math-$46$[2]
@@
The underlying fault in LABEL:listing:lang7 is not the conditional, but the return-statement; removing only the return would also be a valid fix. We instead classify this as an unconditional jump fault. The unwraps-from classification in LABEL:listing:math46 is biased by the fix, when the fault is a definition and use fault (referencing and returning INF instead of NaN). Figure 0(c) shows the percentage correlation between our assigned fault classes with the repair-based classes [2], and Figure 0(d) shows the reverse. We see several one-sided agreements between the classifications, with (cond block, guard) and (expr fix, pred) the only pairs to show two-sided majority agreement. This illustrates that the two classifications are distinct. Figures 0(a) and 0(d) show that both fault and repair class assignments often correlate with the definition fault class, implying that these faults and fixes often focus on definitions.
(a) Fault vs. fault
(b) Repair vs. repair
(c) Repair vs. fault (by fault)
(d) Repair vs. fault (by repair)
@@
Figure 1: Agreement matrices between classes (our fault classes, described in §II; repair classes [2])
V Dataset Usage and Future Research Questions
Our fault classification scheme is useful for its exploration of intrinsic fault characteristics, instead of repairs. We suggest the use of our scheme and classified faults to investigate datasets similar to Defects4J and to reduce debugging costs.
@@ V Dataset Usage and Future Research Questions
We have already considered the research question “Are spectrum-based fault localization (SBFL) techniques more effective for certain fault types?” Our preliminary results show that SBFL performs worse on pure control-flow faults, likely because such faults often involve missing code, which is difficult to localize. SBFL techniques suitable to specific fault types may be developed, then tested on our classified faults and compared with existing techniques. We can also investigate the research question, “Can the classification of an unknown fault be predicted from SBFL performance?”
@@
Other possible applications include automated program repair. For example, we can use the fault class frequency information to guide automated program repair to attempt fixes in order of fault class frequency; if we can classify individual faults on-the-fly, we could even tailor individual repair operations to the specific fault classes.
VI Related Datasets
@@
Goodenough and Gerhart [10] give phase-based error classification (construction, specification, design, requirements), but this is too abstract for our purposes. They also classify program errors into missing control-flow paths, inappropriate path selection, and inappropriate or missing action. These classes are similar to some of our fault classes, but the classification does not account for all faults found in practice.
Duncan and Robson [11] focus on faults in C programs. Their main classes are problematic construct, multiple file usage, and portability issues. However, some of the faults are overly specific (e.g., confusion of & and &&) or not applicable to our dataset (e.g., scanf with wrong types).
@@
Pan et al. [4] describe a classification scheme of 27 bug fix patterns. The scheme is quite fine-grained, e.g., change of if condition expression and change of loop predicate form different classes, despite both being predicate faults. However, since certain fault characteristics can be deduced from the fix patterns, the schema has found use and adoption. Nath et al. [3] use these classes as a fault classification scheme for Java, while Lucia et al. [5] limit their scheme to eight categories, six of which are taken from Pan et al. [4], with two categories added (incl. others). However, predicate faults remain divided between conditionals and loops.
@@
DiGiuseppe and Jones [1] investigate fault localization performance wrt. four fault taxonomies, of which three are quite coarse, with only between four and five fault types. The single fine-grained taxonomy, taken from Hayes et al. [12], is a generic fault taxonomy and includes classes such as visual stimulation and documentation, which are typically not useful in bug datasets. DiGiuseppe and Jones find no correlation between their classes and localization effectiveness; however, our preliminary results show that such a correlation likely exists for our classification, indicating it is more suitable for such applications.
VII Conclusion

### v2

@@ 4. Dataset Statistics
Table 1 shows the frequency of each fault class, in terms of the number of faults it is assigned to, over the 488 classified Defects4J faults. We see that definition fault (data-flow) is the most common fault class for each project. On the other hand, unconditional statement order fault (control-flow) is the least common class. Overall, control-flow classifications outweigh data-flow classifications roughly by a factor of 1.5 to 1, even if the individual control-flow classes are less frequently assigned. Because the fault classes are not disjoint, we consider disjoint fault types, where a fault is assigned only control-flow classes (pure CF), only data-flow classes (pure DF), or both (CF/DF). Pure CF faults outweigh pure DF faults roughly by a factor of 2 to 1, but the majority of faults display a mixture of control- and data-flow characteristics.
@@
Figure 2(a) shows the details of the co-occurrences between our fault classes over the full dataset. Each cell $i,j$ shows the percentage of faults assigned fault class $i$ that are also assigned fault class $j$. There is high co-occurrence between several classes, especially with definition fault. Due to the discussed distribution, faults are often assigned more than one class, and definition fault is the most frequent class.
@@
Sobreira et al. (Sobreira et al., 2018) investigate 395 of these faults; however, they focus on the repair actions and patterns, rather than fault types. The repair actions are very fine-grained, since they distinguish between addition, removal, and modification in nearly all syntactic categories. This leads to 28 different classes. The nine repair patterns are more abstract, and thus closer to our classification. For example, the wraps-with/unwraps-from pattern is comparable with our predicate existence fault class (2.2.2). However, some repair pattern categories are not useful for our purposes. For example, single line does not give much information about the fault itself—the fault is fixed in one line, but this is not the only possible fix. Figure 2(b) shows the co-occurrences of seven repair classes from (Sobreira et al., 2018) over all projects. These generally show lower co-occurrence than our own.
@@
Figure 2(c) shows the percentage correlation between our assigned fault classes with the repair pattern classes (Sobreira et al., 2018), and Figure 2(d) shows the reverse. We see several one-sided co-occurrences between the classifications, with (cond block, guard) and (expr fix, pred) the only pairs to show two-sided majority co-occurrence. The generally low co-occurrence illustrates that the two classifications are distinct. Figures 2(a) and 2(d) show that both fault and repair class assignments often correlate with the definition fault class, implying that these faults and fixes often focus on definitions.
(a) Fault class vs. fault class
(b) Repair class vs. repair class
(c) Repair class vs. fault class (of fault class)
(d) Repair class vs. fault class (of repair class)
@@
Figure 2. Co-occurrence matrices between classes (our fault classes, described in §2; repair classes (Sobreira et al., 2018));
two classes co-occur for a fault if and only if both classes are assigned to the fault
5. Data Availability
The classification results are available at https://doi.org/10.5281/zenodo.15032242. The replication package is available at https://doi.org/10.6084/m9.figshare.28855634.
6. Dataset Usage
Our fault classification scheme is novel for its exploration of intrinsic fault characteristics, not just repairs. Studies that use our classified dataset as a benchmark can thus investigate the performance of novel debugging techniques in terms of the true nature of the faults. The fault classes are generalizable to any programming language in which programs can be represented as control- and data-flow graphs, e.g., Python, and we plan to use our fault taxonomy to investigate the BugsInPy dataset (Widyasari and contributors, 2020).
@@ 6. Dataset Usage
We have already considered the research question “Are spectrum-based fault localization (SBFL) techniques more effective for certain fault types?” Our preliminary results show that SBFL performs worse on pure control-flow faults, likely because such faults often involve missing code, which is difficult to localize. This is supported by Szatmári et al. (Szatmári et al., 2020)’s investigation of the effect of fault fix types on SBFL efficiency: they find that adding an else-branch (predicate block fault—control-flow) is significantly harder to localize, while changing a method call parameter (definition fault—data-flow) is significantly easier to localize than other fix types.
Our classification can be used as a benchmark dataset to validate SBFL techniques that are optimized for specific control- or data-flow fault types, or as a training set for a large language model (LLM) classifier. We can also investigate the research question, “Can the classification of an unknown fault be predicted from SBFL performance?” In other words, if SBFL can localize a fault in the top-$N$ of the ranking, is it more likely to be a data-flow fault?
@@
Other possible applications include automated program repair. For example, we can use the fault class frequency information to guide automated program repair to attempt fixes in order of fault class frequency; we could even tailor individual repair operations to the specific fault classes.
7. Related Work
@@
Goodenough and Gerhart (Goodenough and Gerhart, 1975) give a phase-based error classification (construction, specification, design, requirements), but this is too abstract for any detailed code analysis. They also classify program errors into missing control-flow paths, inappropriate path selection, and inappropriate or missing action. These classes are similar to some of our fault classes, but the classification does not account for all faults found in practice. Duncan and Robson (Duncan and Robson, 1996) focus on faults in C programs. Their main classes are problematic construct, multiple file usage, and portability issues. However, some of the faults are overly specific (e.g., confusion of & and &&) or not applicable to our dataset (e.g., scanf with wrong types).
@@
Pan et al. (Pan et al., 2009) describe a classification scheme of 27 bug fix patterns. The scheme is quite fine-grained, e.g., change of if condition expression and change of loop predicate form different classes, despite both being predicate faults. However, since certain fault characteristics can be deduced from the fix patterns, the schema has found use and adoption. Nath et al. (Nath et al., 2012) use these classes as a fault classification scheme for Java, while Lucia et al. (Lucia et al., 2014) limit their scheme to eight categories, six of which are taken from Pan et al. (Pan et al., 2009), with two categories added (incl. others). However, predicate faults remain divided between conditionals and loops. Szatmári et al. (Szatmári et al., 2020) investigate the effect of these bug fix patterns (specifically if- and sequence-related) on SBFL efficiency. They show that certain patterns are significantly easier (e.g., change to an if condition) or significantly harder (e.g., adding/removing an else branch) to localize than others. Thus localization efficiency can vary within a broader fix pattern, e.g., if-related fixes.
@@
DiGiuseppe and Jones (DiGiuseppe and Jones, 2015) investigate fault localization performance wrt. four fault taxonomies, of which three are quite coarse, with only between four and five fault types. The single fine-grained taxonomy, taken from Hayes et al. (Hayes et al., 2011), is a generic fault taxonomy and includes classes such as visual stimulation and documentation, which are typically not useful in bug datasets. DiGiuseppe and Jones find no correlation between their classes and localization effectiveness; however, our preliminary results show that such a correlation likely exists for our classification, indicating it is more suitable for such applications. Catolino et al. (Catolino et al., 2019) analyze the bug reports of $119$ GitHub projects and form a taxonomy of nine high-level fault classes: configuration, network, database-related, GUI-related, performance, permission/deprecation, security, program anomaly, test code-related. Humbatova et al. (Humbatova et al., 2020) investigate the classification of deep learning faults, forming a taxonomy of five top-level categories (model, tensors & input, training, GPU usage, API), with model, tensors & input, training each having multiple subcategories.
In recent years, researchers have investigated the use of deep learning models in automated software fault classification. Li et al. (Li et al., 2024) use a pre-trained BERT model to categorize faults, and show that their model improves on existing models. However, their fault classes (hardware, algorithm, software, human factors, interaction, network) are rather abstract and unsuited for fault localization and repair benchmark datasets, which generally consist of software faults. Furthermore, the automated fault classification was tested on “detailed case descriptions” (Li et al., 2024), which are not always available in fault datasets, and require a programmer’s understanding of the fault in question. This type of classification may not capture the full nuance of the fault.
8. Conclusion and Future Work
@@
In this paper we report intermediate results from ongoing work. Current work includes a more expansive classification, performed by a larger group of raters on larger-scale datasets and different programming languages (e.g., BugsInPy (Widyasari and contributors, 2020)), using a more detailed categorization. We also plan to consider dynamic flow graphs and to test scalability and statistical significance among the fault classes. We can further investigate the effect of different fault classes on fault localization and repair ability. Finally, automated analysis of the graph structure may be possible, but the use of LLMs or deep neural networks (DNNs) for automatic fault classification and prediction is also worth considering.